1.) The following relations describe monthly demand and supply for a wheat

2 3р QD 32 Р 4 ЕР- 16 P

where P is the price (in cents) per pound and Q is the quantity (in millions) of pounds.

Suppose government passes a regulation which makes it illegal for buyers to buy at a price lower than 60 cents per pound.

a.) Calculate the value of consumer surplus post-regulation. Compare it with pre-regulation consumer surplus.

b.) Calculate the value of producer surplus post-regulation. Compare it with pre-regulation producer surplus.

c.) Identify other undesirable effects of this regulation.
Answer #1

Em Op=Os 32 - 2 P = 4 P-16 48 = 8 p=48x5 40 oly = 32- 32- 165 16 8 16 At luce = 60, do -32-2 160) = 32-24=8 Consumer surples